In the system shown, a horizontal force Fx = 40 N acts on an object of mass m2 = 8.00 kg. A friction force fk = 5 N acts between object m2 and the table top, and m1 = 2.00 kg. i) Calculate the acceleration of the masses. ii) Calculate the tension in the string.
1
Expert's answer
2020-05-28T13:23:39-0400

"2\; Newton's\; law:\\1)\;F_{fr2}=\mu\times m_2\times g;\mu=\frac{F_{fr}}{m_2\times g}=\frac{5}{8\times 9.8}=0.064\\\\(m_1+m_2)\times a=F_x-F_{fr1}-F_{fr2}\\a=\frac{40-0.064\times 2\times 9.8-5}{10}=3.375\frac{m}{s^2};\\2)\;m_1\times a=T-F_{fr1};T=m_1\times a+F_{fr1}\\T=2\times3.375+0.064\times 2\times 9.8=\\=6.75+1.25=8N;\\Answer:\;the\; acceleration\; of\; the\; masses:\\3.375\frac{m}{s^2};the \;tension\; in\; the\; string:8N"